OpenStudy (anonymous):

Consider the graph of Bx + Cy = A where A, B, and C are all positive constants. (a) Find the coordinates of the x-intercept. Use upper case letters. Find the coordinates of the y-intercept. Use upper case letters. (b) Is the graph increasing or decreasing?

OpenStudy (science0229):

That equation would be y=-(B/C)x +A... x-intercept is when y=0. So, x=A/B And, -B/C is negative, so the graph is decreasing... Final Answer a) x-intercept=A/B b) The graph is decreasing

OpenStudy (anonymous):

thanks i got it :)